Tutor could you please help me with the following problem: x+4y=16
If both x and y are positive integers in this equation, how many different values can x have? (1,2, 3 or 4) I chose 3..am I correct?
Thank you for your help!

Answer by checkley71(8403)

X+4Y=16
3+4Y=16
4Y=16-3
4Y=13
Y=13/4 WRONG ANSWER FOR Y.
------------------------------------------
4+4Y=16
4Y=16-4
4Y=12
Y=3 ONLY ONE THAT WORKS.